# Periodic orbits and gravitational wave radiation in short hair black hole spacetimes for an extreme mass ratio system

### v2

@@ VI Discussion and conclusions
In this paper, the short hair BH model we study is static and spherically symmetric, although its physical background is similar to the real-universe environment (that is, the case of considering anisotropic matter instead of vacuum). However, it is well-known that BHs in the real universe usually have spins. Therefore, although the static BH model may be a good approximation in the region far from the BH, when studying the particle motion close to the BH’s event horizon, the introduction of spin effects may reveal more interesting physical phenomena. In particular, the frame-dragging effect caused by spin may enhance the quantum effect and significantly affect the gravitational wave signal or the orbital behavior of test particles, which may lead to novel observational results. Therefore, considering spin effects and reconstructing more accurate gravitational wave waveforms are of great significance for in-depth understanding of the physical properties of the short hair BH. In future research, we plan to further explore the complex nature of the short hair BH by analyzing the timelike geodesic dynamics in a rotating short hair BH and constructing accurate gravitational wave waveforms.

### v3

@@ VI Discussion and conclusions
In this paper, the short hair BH model we study is static and spherically symmetric, although its physical background is similar to the real-universe environment (that is, the case of considering anisotropic matter instead of vacuum). However, it is well-known that BHs in the real universe usually have spins. Therefore, although the static BH model may be a good approximation in the region far from the BH, when studying the particle motion close to the BH’s event horizon, the introduction of spin effects may reveal more interesting physical phenomena. In particular, the frame-dragging effect caused by spin may enhance the quantum effect and significantly affect the gravitational wave signal or the orbital behavior of test particles, which may lead to novel observational results. Therefore, considering spin effects and reconstructing more accurate gravitational wave waveforms are of great significance for in-depth understanding of the physical properties of the short hair BH. In future research, we plan to further explore the complex property of the short hair BH by analyzing the timelike geodesic dynamics in the rotating short hair BH and constructing precise gravitational wave waveforms (such as taking high-order multipole moments and non-adiabatic situations into consideration).